Grade 12th pass9 grade science

Write the names of the following compounds and deduce their molar masses. Na2​SO4​,K2​CO3​,CO2​,MgCl2​,NaOH,AlPO4​,NaHCO3​

Harshit Singh
4 Years agoGrade 12th pass

1 Answer

Jayesh Vashishtha
4 Years ago
Dear student,
 
Na2​SO4​ : Sodium sulphate 
Molecular Mass : 23(2)+32+16(4) = 46+32+64 = 142 g/mol
 
K2​CO3​ : Potassium carbonate
Molecular Mass : 39(2)+12+16(3) = 78+12+48 = 138 g/mol
 
CO2​ : Carbon dioxide
Molecular Mass : 12+16(2) = 12+32 = 44 g/mol
 
MgCl2​ : Magnesium chloride
Molecular Mass : 24.31+35.5(2) = 95.31 g/mol
 
NaOH : Sodium hydroxide
Molecular Mass : 23+16+1 = 40 g/mol
 
AlPO4​ : Aluminium phosphate
Molecular Mass : 27+31+16(4) = 27+31+64 = 122 g/mol
 
NaHCO3​ : Sodium bicarbonate
Molecular Mass : 23+1+12+16(3) = 23+1+12+48 = 84 g/mol
 
 
regards,
Team Askiitians